OpenStudy (anonymous):

order the following from least to greatest. √2, -2/3, √11, 0.8, -0.2 a. √2, -0.2, -2/3, √11, 0.8 b. -2/3, -0.2, √2, √11, 0.8 c. -2/3, -0.2, 0.8, √2, √11 d. √2, √11, 0.2, -2/3, 5.1

OpenStudy (igreen):

\(\sqrt2 \approx 1.41\) \(-\dfrac{2}{3} \approx -0.67\) \(\sqrt{11} \approx 3.32\) \(0.8\) \(-0.2\) Can you order these from least to greatest now? @taytay105

OpenStudy (igreen):

It's going to be: \(-\dfrac{2}{3}, -0.2, 0.8, \sqrt2, \sqrt{11}\)

OpenStudy (igreen):

Therefore, your answer is C.

OpenStudy (anonymous):

C?

OpenStudy (igreen):

Yes.